OpenStudy (anonymous):

Simplify x^2-12x+20/x-10 @satellite73 @phi

OpenStudy (anonymous):

x-2 x-10/ x-10= x-2

OpenStudy (phi):

you have to factor the top the +20 (3rd #) means write down pairs that multiply to give you 20 1,20 2,10 4,5 the + in +20 means that both numbers of the pair have the same sign... either both + or both - -12 from -12x, the minus sign tells you the largest of the pair is negative, (since both have the same sign ... both are - and they two combine to give you -12 now look at our pairs: make them minus, and add up: -1 + -20 = -21 -2 + -10 = -12 -4 + -5 = -9 the second pair give you the -12, so that is the pair to use as factors: (x-2)(x-10)

OpenStudy (phi):

so the problem is (x-2)(x-10)/(x-10) the (x-10)/(x-10) is 1 so you get (x-2)